Why is the change in pH of 3.00 mL of 0.100M pure acetic acid (weak acid that does not ionize fully) greater than the change in pH of a 3.00 mL sample of a .100 M acetic acid + 0.500-gram acetate buffer solution when you add 2 drops (1x10^-4 L) of 3M NaOH? Please show with equations and calculations!
Answer #1

Answer – We are given, [CH3COOH] = 0.100 M, volume = 3 mL

Acetic acid is weak acid and it does not dissociate completely, so we need to put ICE

    CH3COOH + H2O ------> H3O+ + CH3COO-

I    0.100 0               0

C    -x                                 +x            +x

E 0.100-x                         +x             +x

Ka = [H3O+] [CH3COO-] / [CH3COOH]

1.8*10-5 = x*x /(0.100-x)

1.8*10-5 *(0.100-x) = x2

Now we know the Ka value is too small, so we can neglect x in the 0.100-x

1.8*10-5*0.100 = x2

x = 0.00134M

x = [H3O+] = 0.00134 M

pH = -log [H3O+]

      = -log 0.00134 M

       = 2.87

Now we need to calculate change in pH after added 2 drops (1x10^-4 L) of 3M NaOH

Moles of acetic acid = 0.100 M * 0.003 L

                                    = 0.0003 moles

Moles of NaOH = 3 M * 1*10-4 L

                         = 0.0003 moles

Reaction –

    CH3COOH + NaOH ------> H2O + CH3COONa

      0.0003     0.0003                          0.0003

so mole of both acid and base are equal, so there is formation of conjugate baase only

moles of CH3COO- = 0.0003 moles

total volume = 3 mL + 1.0*10-7 mL = 3 mL

[CH3COO-] = 0.0003 moles / 0.003 L

                    = 0.100 M

Now ICE chart for conjugate base

    CH3COO- + H2O ------> HO- + CH3COOH

I    0.100 0               0

C    -x                              +x            +x

E 0.100-x                         +x             +x

Kb = [HO-] [CH3COOH] / [CH3COO-]

Kb = 1*10-14 / 1.8*10-5 = 5.56*10-10

5.56*10-10= x*x /(0.100-x)

5.56*10-10*(0.100-x) = x2

Now we know the Kb value is too small, so we can neglect x in the 0.100-x

5.56*10-10*0.100 = x2

x = 7.45*10-6 M

x = [HO-] = 7.45*10-6 M

pOH = -log [HO-]

      = -log 7.45*10-6 M

       = 5.12

pH = 14 – 5.12

      = 8.87

so pH change = 8.87 – 2.87

                          = 6

so pH difference is 6

Now we need to calculate pH difference for the buffer solution –

3.00 mL sample of a .100 M acetic acid + 0.500-gram acetate buffer solution

Moles of acetate CH3COO- = 0.500 g / 59.045 g.mol-1

= 0.00847 moles

[CH3COO-] = 0.00847 moles / 0.003 L

                   = 2.82 M

Now using the Henderson Hasselbalch equation

pH = pKa + log [CH3COO-] / [CH3COOH]

      = 4.74 + log 2.82 M /0.100 M

      = 4.74 + 1.45

       = 6.19

Now pH after added NaOH

Moles of acetic acid = 0.100 M * 0.003 L

                               = 0.0003 moles

Moles of NaOH = 3 M * 1*10-4 L

                         = 0.0003 moles

Moles of CH3COO- = 0.00847 moles

When we added NaOH moles of base increase and moles of acid decreased

Moles of acetic acid = 0.0003 – 0.0003 moles = 0

Moles of CH3COO- = 0.00847 moles + 0.0003 = 0.00877 moles

total volume = 3 mL + 1.0*10-7 mL = 3 mL

[CH3COO-] = 0.00877 moles / 0.003 L

                    = 2.92 M

Now ICE chart for conjugate base

    CH3COO- + H2O ------> HO- + CH3COOH

I    2.92 0               0

C    -x                             +x            +x

E 2.92-x                         +x             +x

Kb = [HO-] [CH3COOH] / [CH3COO-]

Kb = 1*10-14 / 1.8*10-5 = 5.56*10-10

5.56*10-10= x*x /(2.92-x)

5.56*10-10*(2.92-x) = x2

Now we know the Kb value is too small, so we can neglect x in the 2.92-x

5.56*10-10*2.92 = x2

x = 4.02*10-5 M

x = [HO-] = 4.02*10-5 M

pOH = -log [HO-]

      = -log 4.02*10-5 M

       = 4.39

pH = 14 – 4.39

      = 9.60

so pH change = 9.60 – 6.19 = 3.41

so pH change of 3.00 mL of 0.100M pure acetic acid is 6 and pH change of a 3.00 mL sample of a .100 M acetic acid + 0.500-gram acetate buffer solution is 3.41. So 3.00 mL of 0.100M pure acetic acid greater than the change in pH of a 3.00 mL sample of a .100 M acetic acid + 0.500-gram acetate buffer solution.